Q2 The following data relate to a manufacturing department for a period: Budgeted data for the...

Q2

The following data relate to a manufacturing department for a period:

Budgeted data for the coming year are as follows:

Direct labour hours 60,000 hours

Machine hours 55,000 hours

Direct labour cost $110,000

Direct material cost $125,000

Production overhead $70,400

Actual data for the period are as follows:

Direct labour hours 55,500 hours

Machine hours 40,000

Direct labour cost $125,000

Direct material cost $150,000

Production overhead $67,800

Required:

  1. Calculate the production overhead absorption rate predetermined for the period based on :
  1. Machine hours;
  2. Direct labour hours
  1. Two pieces of furniture are to be manufactured for customers. Direct costs are as follows:

Job 123                                             Job 124

Direct material    $15,400                                             $10,800

Direct labour        210 hours Dept. A         @$16/ hour:         160 hours Dept. A @20/hour

                             120 hours Dept. B @$12/hour; 100 hours Dept. B @$18/hour

                             100 hours Dept. C @$14/hour; 140 hours Dept. C@$10/hour

Selling and administrative overheads for each job are 10% of factory cost.

Required:

(b)     Calculate the total costs of each job using all the two bases in question (1) above.

(c) If the firm quotes prices to customers that reflect a required profit of 25 per cent on selling price, calculate the quoted selling price for each job.
Answer #1

Production overhead absorption rate predetermined for the period based on :

.i) Machine hours:

Budgeted Production overhead= $70,400

Budgeted Machine hours =55,000 hours

Predetermined Production overhead absorption rate=70400/55000=$1.28

Ii)Direct labour hours

Budgeted Production overhead= $70,400

Budgeted Direct labor hours =60,000 hours

Predetermined Production overhead absorption rate=70400/60000=$1.17

Calculate the total costs of each job using all the two bases:

JOB 123

Direct material    $15,400

Direct Labor: Dept A=210*$16=$3360

Direct Labor Dept B=120*$12=$1440

Direct Labor Dept C=100*$14=$1400

Total Direct Labor=3360+1440+1400=$6200

Total hours=210+120+100=430

Production Overhead=$1.17*430=$505

Total Factory Cost of Job 123=15400+6200+505=$22,105

Selling and Administrative overhead=10%*22105=$2211

TOTAL COST OF Job 123=22105+2211=$24,316

JOB 124

Direct material    $10,800

Direct Labor: Dept A=160*$20=$3200

Direct Labor Dept B=100*$18=$1800

Direct Labor Dept C=140*$10=$1400

Total Direct Labor=3200+1800+1400=$6400

Total hours=160+100+140=400

Production Overhead=$1.17*400=$469

Total Factory Cost of Job 124=10800+6400+469=$17,669

Selling and Administrative overhead=10%*17669=$1767

TOTAL COST OF Job 124=17669+1767=$19,436

(c) If the firm quotes prices to customers that reflect a required profit of 25 per cent on selling price, calculate the quoted selling price for each job.

Assume Quoted Price =P

Profit=0.25P

Cost=P-0.25P=0.75P

Quoted Price=P=Cost/0.75

Quoted Price of Job 123=24316/0.75=$32,421

Quoted Price of Job 124=19436/0.75=$25,915